# Virtual Reflections on a Dynamic 2D Eye Model Improve Spatial Reference Identification

### v3

@@ 4.4 References in physical space
Looking back at the original objective of the present investigation, i.e., the improvement of spatial reference communication with 2D eye models, a new direction for follow-up research can arise from an application of the approach within more physically instantiated machine interaction partners, such as robots that possess a movable head.
Previous work has shown that humans perceive the direct gaze of a robot more clearly when the robot’s eye gaze is coordinated with its head movement, as opposed to when the head remains stationary [2]. In the present study, all three eye conditions dynamically tracked the user’s face, and the results demonstrate that participants could accurately judge the machine mediator’s gaze direction, despite its static screen-based mock-up head. Previous research also suggests that combining eye and head movements while simulating visual perception processing enhances the perception of a robot’s attentional gaze [2]. Moving forward, we aim to incorporate the mirror-eye functionality into a physical robot to further examine the expressive potential that this approach may afford in coordination with physical action.
@@ 4.5 Conclusion
The introduction of a reflection-like overlay on the display of virtual eye models (mirror eyes) significantly improved face identification accuracy compared to a pure, non-reflective eye model without measurably affecting identification speed. In congruence with this accuracy gain, the subjective user experience in a game-like interaction with the eye models significantly improved with the mirror eyes compared to conditions that either lacked the reflection capability or the eye-like appearance. These results suggest that the benefits associated with mirror eyes cannot be attributed to the underlying eye model or the reflection alone, but only to their combination. Subsequent research should investigate the transfer of such effects to the reflections of other classes of gaze targets as well as extended applications in more physical embodiments.

### v4

@@ 5.5 References in physical space
Looking back at the original objective of the present investigation, i.e., the improvement of spatial reference communication with 2D eye models, one aspect that requires closer examination is depth of gaze. Although our eye model encodes focus distance in the interpupillary distance and reflections naturally become smaller with increasing distance, the present study did not investigate the effects of Mirror Eyes at different levels of focus depth.
Another direction for follow-up research can arise from an application of the approach within more physically instantiated machine interaction partners, such as robots that possess a movable head.
Previous work has shown that humans perceive the direct gaze of a robot more clearly when the robot’s eye gaze is coordinated with its head movement, as opposed to when the head remains stationary [2, 26]. In the present study, all three eye conditions dynamically tracked the user’s face, and the results demonstrate that participants could accurately judge the machine mediator’s gaze direction, despite its static screen-based mock-up head. Previous research also suggests that combining eye and head movements while simulating visual perception processing enhances the perception of a robot’s attentional gaze [2]. Also, potential cross-modal interactions with other output modalities such as speech (e.g.,  [62]) should be considered in future developments.
Moving forward, we aim to incorporate the Mirror Eye functionality into a physical robot to further examine the expressive potential that this approach may afford in coordination with physical action.
@@ 5.6 Conclusion
The introduction of a reflection-like overlay on the display of virtual eye models (Mirror Eyes) significantly improved reference identification accuracy compared to a pure, non-reflective eye model without affecting identification speed. In congruence with this accuracy gain, the subjective user experience in a game-like interaction with the eye models significantly improved with the Mirror Eyes compared to conditions that either lacked the reflection capability or the eye-like appearance. These results suggest that the benefits associated with Mirror Eyes cannot be attributed to the underlying eye model or the reflection alone, but only to their combination. Subsequent research should investigate the transfer of such effects to the reflections of other classes of gaze targets, explore options for a selective modulation of mirroring for attention guidance, and ascertain meaningful coordination with degrees of freedom available to more physically integrated machine agents.
Acknowledgments
We thank Christian Arzate, Yotam Sechayk, and Heiko Wersing for their thoughtful discussions, which were instrumental in shaping the user study design.